A project has the following estimated data: price = $85 per unit; variable costs = $46.75 per unit; fixed costs = $5,600; required return = 14 percent; initial investment = $10,000; life = three years. Ignore the effect of taxes.

What is the accounting break-even quantity?

What is the cash break-even quantity?

What is the financial break-even quantity?

What is the degree of operating leverage at the financial break-even level of output?
Answer #1

1.Accounting break even quantity=Fixed cost/(sales price per unit-variable cost per unit)=5600/(85-46.75)=146.41 units

2.Cash break even quantity=(Fixed cost-depreciation)/(sales price per unit-variable cost per unit)=(5600-3333.33)-(85-46.75)=59.26 units

-->Depreciation=investment/years=10000/3=3333.33

3.Finance break even quantity=(Fixed cost+required return)/(sales price per unit-variable cost per unit)=(5600+1400)/(85-46.75)=183.01 units

-->Required return=Initial investment*required return=10000*14%=1400

4.Degree of Operating leverage at financial break even output=(Sales-variable cost)/(Sales-variable cost-fixed cost)=(15555-8555.25)/(15555-8555.25-5600)=5

Particulars 183 Units
Sales price 15555
Variable cost 8555.25
Contribution 6999.75
Fixed cost 5600